OpenStudy (anonymous):

how do i even start this? Paul invested twice as much money in an account paying 4% interest than he did in an account paying 2% interest. If the total interest paid was $250, how much did he invest in each?

terenzreignz (terenzreignz):

You start by representing the unknown values... for instance We can let x = how much Paul invested in the 2% interest account. If that were the case, how do you represent how much Paul invested in the 4% account? :)

OpenStudy (anonymous):

2x+.04

terenzreignz (terenzreignz):

no, just 2x. It was specified that Paul invested TWICE as much in that particular account, so it's 2x. Get it? :)

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

Ok. So the interest gained in the 2% account is...?

OpenStudy (anonymous):

125

terenzreignz (terenzreignz):

what? How do you figure that?

OpenStudy (anonymous):

idk how to figure it out

terenzreignz (terenzreignz):

It's 2% of x (which is how much he deposited in that account)

OpenStudy (anonymous):

so x divided by .02

OpenStudy (anonymous):

i'm supposed to solve this by using a system of equations

terenzreignz (terenzreignz):

Well, that changes things... but the essence is the same... We can let x = amount he deposited in the 2% account y = amount he deposited in the 4% account ok? :)

OpenStudy (anonymous):

okay

terenzreignz (terenzreignz):

so, the interest in the 2% account is 2% of x, right? or 0.02x any questions? :)

OpenStudy (anonymous):

nope i understandthat

terenzreignz (terenzreignz):

And the interest in the 4% account is...?

OpenStudy (anonymous):

.04y

terenzreignz (terenzreignz):

ok, so the total interest all in all, which is basically the sum of 0.02x and 0.04y, is...?

OpenStudy (anonymous):

250

terenzreignz (terenzreignz):

so... we have one equation 0.02x + 0.04y = 250 We need another, can you get one? :)

terenzreignz (terenzreignz):

we know that what he invested in the 4% account is TWICE what he deposited in the 2% account. So y = 2x, aye? :)

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

so, you now have two equations, can you solve it from here? :)

OpenStudy (anonymous):

i did it wrong. 0.02x+0.04(2x)=250 0.04x+0.04=250

OpenStudy (anonymous):

ohhhh no

terenzreignz (terenzreignz):

You forgot to retain the x... look again :)

OpenStudy (anonymous):

so it's 0.04x+0.08x=250

terenzreignz (terenzreignz):

that's right. Now figure it :)

terenzreignz (terenzreignz):

Wait NO

terenzreignz (terenzreignz):

Why did the 0.02x become 0.04x?

OpenStudy (anonymous):

just kidding

OpenStudy (anonymous):

so .1x=250

terenzreignz (terenzreignz):

That's right :) now solve for x

OpenStudy (anonymous):

2500

terenzreignz (terenzreignz):

Nicely done. Now, with y = 2x, what is y?

OpenStudy (anonymous):

5000

terenzreignz (terenzreignz):

And you now have your values :)

OpenStudy (anonymous):

THANkyou! can you help me with another problem you're really good at explaining things

terenzreignz (terenzreignz):

Sure, hit me :)

OpenStudy (anonymous):

Chuck and Dana agree to meet in chicago for the weekend. Chuck travels 220 miles in the same time that Dana travels 205 miles. if Chuck's rate of travel is 3mph more than Dana's then at what rate does chuck travel?

terenzreignz (terenzreignz):

Are you to use systems again?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

Hang on a sec

terenzreignz (terenzreignz):

I can solve it, but not with systems, just a basic linear equation in one variable.

terenzreignz (terenzreignz):

Well, actually, no, It won't be linear.

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

i guess still show me?

terenzreignz (terenzreignz):

Hang on, I was wrong, sorry

terenzreignz (terenzreignz):

All right, systems is the way to go here, ready? This might be tricky...

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

First, we let x = Chuck's rate and we let y = Dana's rate, ok?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

We know that Chuck's rate is 3mph more than Dana's, so x = y + 3 any questions? :)

OpenStudy (anonymous):

nope

terenzreignz (terenzreignz):

Ok, we have one equation, we need another, which will be a bit tricky, so pay attention. You know that speed = distance over time, right?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

So time = distance over speed , right?

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

so, they arrive at the same time, so time is constant. The time = Chuck's distance / Chuck's speed and also time = Dana's distance / Dana's speed are you with me so far?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

So Chuck's distance / Chuck's speed = Dana's distance / Dana's speed right?

OpenStudy (anonymous):

yeah

terenzreignz (terenzreignz):

What distance did Chuck and Dana travel? 220 and 205, right? And Chuck's speed is x, Dana's speed is y so you get \[\frac{220}{x}=\frac{205}{y}\]

terenzreignz (terenzreignz):

There's your second equation, so you have your system, it shouldn't be hard to evaluate Chuck's speed (x) :D Have fun

terenzreignz (terenzreignz):

remember x = y + 3

OpenStudy (anonymous):

yes

terenzreignz (terenzreignz):

so go ahead and solve the systems :)

OpenStudy (anonymous):

i got 41 for y

OpenStudy (anonymous):

and 44 for x

terenzreignz (terenzreignz):

Nicely done :)

OpenStudy (anonymous):

thank you! youre really good at explaining i get it when you explain it

terenzreignz (terenzreignz):

For the record, all you needed was the value for x, which you did get, nicely done, and thanks! :)

OpenStudy (anonymous):

no problem!